OpenStudy (anonymous):

Write a recursive definition for the following sequence. 2, -4, 8, -16

jimthompson5910 (jim_thompson5910):

do you notice any patterns at all?

jimthompson5910 (jim_thompson5910):

if not, then try thinking of each term as positive 2,4,8,16 what's the pattern here?

OpenStudy (anonymous):

x -2?

jimthompson5910 (jim_thompson5910):

times -2, yes

jimthompson5910 (jim_thompson5910):

2 times -2 = -4 -4 times -2 = 8 etc etc

jimthompson5910 (jim_thompson5910):

so the recursive rule is this \[\Large a_{1} = 2\] \[\Large a_{n} = -2*a_{n-1}\] the first line above states the first term is 2. The next line tells you how to generate the other terms (multiply each by -2)

OpenStudy (anonymous):

Thank you,

OpenStudy (anonymous):

*!

jimthompson5910 (jim_thompson5910):

you're welcome